what is d [OH] of a 0.2M solution of ammonia.given,k=1.8*1/100000
1
Expert's answer
2012-11-20T08:49:27-0500
Ammonia is a weak base, so the most convenient approach is to calculate [OH-] using Kb

[OH-] = (K*C)1/2 = (1.8*1/100000*0.2)1/2 = 0.001897 M